OpenStudy (anonymous):

f(x)=(6/x)-2 use a four step process to find f'(x) and find f'(1), f'(2),f'(3). this question has to do with primes.

jimthompson5910 (jim_thompson5910):

f(x) = (6/x) - 2 f(x) = 6x^(-1) - 2 f'(x) = -6x^(-2) --------------- f'(1) = -6(1)^(-2) = -6/(1^2) = -6/1 = -6 f'(2) = -6(2)^(-2) = -6/(2^2) = -6/4 = -3/2 f'(3) = -6(3)^(-2) = -6/(3^2) = -6/9 = -2/3